OpenStudy (anonymous):

@austinL What do I do when I have (-9y+12+9y=12)

OpenStudy (austinl):

\(-9y+12+9y=12\) \(-9y+9y=12-12\) \(0=0\) That should mean that there are infinitely many solutions. I believe.

OpenStudy (anonymous):

Sorry, my computer froze up. Ahh! That's what I thought. So it's an Identity equation.

OpenStudy (austinl):

I suppose, I am a little rusty with the technical jargon.

OpenStudy (anonymous):

Haha, you're a genius is what you are =) Thank you so much! =)